OpenStudy (liliakarina):

What's the distance between -5/7 and -2/5?, I say it's -11/35

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

\[-\frac 57 - (-\frac 25) \implies -\frac 57 + \frac 25 = -\frac{11}{35}\]

OpenStudy (liliakarina):

All right!!